Circular No. 6081

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

PERIODIC COMET MACHHOLZ 2 (1994o)

Improved orbital elements from MPC 23956:

T = 1994 Sept.18.8010 TT  ω = 149.^{o}2566
e = 0.750261  Ω = 246.1808 }2000.0
q = 0.752551 AU  i = 12.7877

a = 3.013353 AU n^{o} = 0.1884209 P = 5.231 years

1994TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Sept. 25 9^{h}06.^{m}69 +20^{o}55.^{′}9 0.731 0.760 49.^{o}0 84.^{o}5  8.0
30 9 20.42 +17 33.2 0.791 0.777 49.7 79.4  8.3
Oct. 5 9 33.52 +14 26.7 0.848 0.802 50.6 74.6  8.7
10 9 45.99 +11 34.4 0.902 0.834 51.7 70.1  9.1
15 9 57.83 + 8 54.8 0.953 0.873 53.1 66.0  9.5
20 10 09.01 + 6 26.4 1.000 0.916 54.7 62.4  9.9
25 10 19.50 + 4 08.1 1.042 0.964 56.4 59.3 10.3
30 10 29.29 + 1 59.0 1.080 1.014 58.4 56.5 10.8
Nov. 4 10 38.37 − 0 01.8 1.113 1.067 60.6 54.1 11.2
9 10 46.70 − 1 54.9 1.141 1.121 63.0 51.9 11.5
14 10 54.29 − 3 40.9 1.165 1.176 65.6 50.0 11.9
19 11 01.11 − 5 20.2 1.185 1.232 68.4 48.2 12.2
24 11 07.16 − 6 53.4 1.200 1.288 71.4 46.6 12.5

The orbital elements and ephemeris above refer to the principal nucleus, designated component A. Components B, C, D and E are designated in order eastward from component A and are thus the fourth, third, second and fifth components, respectively, mentioned on IAUC 6071, 6070, 6066, and 6071. Their positions can be described well with the above orbital elements, provided that T is changed to Sept. 18.866, 19.199, 19.206 and 19.267, respectively.

Total visual magnitude and coma-diameter estimates of component B (cf. IAUC 6071): Sept. 9.45 UT, 12.0, 2′ (R. A. Keen, Mount Thorodin, CO, 0.32-m reflector); 11.51, 11.9, 1′.4 (C. S. Morris, Pine Mountain Club, CA, 0.26-m reflector).

Total visual magnitude and coma-diameter estimates of component D (cf. IAUC 6074): Sept. 9.45 UT, 11.0, 3′ (Keen); 11.48, 11.3, 3′.3 (Morris).

1994 September 21 Daniel W. E. Green